2. i. Define simple harmonic motion

iii. Show that a mass, m, hung on a spring of spring constant, k, when


1
Expert's answer
2021-03-17T16:55:29-0400

In simple harmonic motion, the acceleration of the system, and therefore the net force, is proportional to the displacement and acts in the opposite direction of the displacement.

A good example of SHM is an object with mass m attached to a spring on a frictionless surface The object oscillates around the equilibrium position, and the net force on the object is equal to the force provided by the spring. This force obeys Hooke’s law.


"a+\omega^2x=0\\x=A\cos{(\omega t+\phi)}"

Where


"\omega=\sqrt{\frac{k}{m}}"